A 75 kg bobsled is pushed along a horizontal surface by two athletes. After the bobsled is pushed a distance of 5.0 m starting from rest, its speed is 6.0 m/s. Find the NET FORCE on the bobsled.
1
Expert's answer
2020-03-30T07:52:18-0400

The change of energy is equal to work done

"K_f-K_i=W=Fd."

Hence, the applied net force

"F=\frac{K_f-K_i}{d}=\frac{mv_f^2}{2d}=\frac{75\times 6.0^2}{2\times 5.0}=270\:\rm N."